OpenStudy (anonymous):

The price of 9-volt batteries is increasing according to the function below, where t is years after January 1, 1980. During what year will the price reach $5? P(t) = 1.1 e0.047t A. 2016 B. 2008 C. 2004 D. 2012

OpenStudy (kropot72):

\[\large P(t)=5=1.1e^{0.047t}\] \[\large e^{0.047t}=\frac{5}{1.1}\] Taking natural logs of both sides, we get: \[\large 0.047t=\ln \frac{5}{1.1}=\ln 4.54545\] Therefore the required value of t is given by: \[\large t=\frac{\ln 4.54545}{0.047}=you\ can\ calculate\]

OpenStudy (kropot72):

Don't forget to add the value of t to 1980 to find the year.